OpenStudy (anonymous):

what is the slope of the line through the given point (-3,1) and (5,4)

OpenStudy (anonymous):



OpenStudy (anonymous):

Question? How did you come up with that? I thought the (-3,1) would be in the II quadrant and the (5,4) in the i quadrant. Wouldn't the IV quadrant be a (-1,-3)?

OpenStudy (anonymous):

put it like this X-axis = -3 and 5 and Y-axis = 1 and 4 after you plot those points you line them up

OpenStudy (anonymous):

Oh,wait.. is that the m=y2-y1 thing?

OpenStudy (anonymous):

yes sir

OpenStudy (anonymous):

Mame..lol.. Im so excited, I put two and two together.. thank you!

OpenStudy (anonymous):

your welcome........mame :)

OpenStudy (anonymous):

So is the slope of the line 3/2?

OpenStudy (anonymous):

wait what? no that's it cause your question says that you only the slope unless there's more you didn't say?

OpenStudy (anonymous):

The question is: Find the slope of the line through the given points (-3,1) and (5,4)

OpenStudy (anonymous):

yeah what i did is the answer that's all you need to get it right :)

OpenStudy (anonymous):

slope m= (y2-y1)/(x2-x1)

OpenStudy (anonymous):

Ok, Mark so what I think your saying is that it is: M=4-1 / 5-3 ~ M=3/2

OpenStudy (anonymous):

almost there (4-1)/(5-(-3))+....?

OpenStudy (anonymous):

My logic says it has to equal 8, but I have no idea how to get it to equal 8

OpenStudy (anonymous):

or m=(1-4)/(-3-5)=....?

OpenStudy (anonymous):

m=(1-4)/(-3-5)=.3/8...? or =(4-1)/(5-(-3))=+3/8....? yes or no?

OpenStudy (anonymous):

AHHHHHHH... I GOT IT !!!!! WHOOT WHOOT

OpenStudy (anonymous):

yeheeeyyyyyyyy .... :D good luck now have fun :D

OpenStudy (anonymous):

The night is still young and I have 4 'sections' to get through.. lol

OpenStudy (anonymous):

just dont forget to give Bweezy5 a medal.... thanks :D

OpenStudy (anonymous):

BWEEZY5 ROCKS and SO DO YOU!!! How do I give medals???

OpenStudy (anonymous):

wait VAT!!! lol i didn't understand anything guys were doin O.o

OpenStudy (anonymous):

formula for slope \[slope m=\frac{ \Delta y }{ \Delta x } = \frac{ y _{2}- y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (anonymous):

I have another slope and Y intercept question. Do I close out here and post a new question?

OpenStudy (anonymous):

jesus is that all it says???

OpenStudy (anonymous):

yes mame close this one and say the next question

OpenStudy (anonymous):

use the formula at. given point (-3,1)=(x1,y1) and (5,4)=(x2,y2)

OpenStudy (anonymous):

no wait lol you for got to close it

OpenStudy (anonymous):

@ ChristaB to give medal just push the buytton there at the further right..lol :D